# Local EGOP for Continuous Index Learning

### v2

@@ B.3.5 Matrix Recurrence
For this to occur, it must be that both $X_{2},X_{1}$ have eigenvalue $e^{-c}$ with a common eigenvector $v$. As the sequence is rational in $X_{2},X_{1}$, $v$ will be an eigenvector for all subsequent matrices $X_{i}$. Notice $\log\lambda_{v}(X_{4})=-\log(\eta e^{c}+(1-\eta)e^{-c})\in(-c,c)$, hence $\sup_{i>0}|\log\lambda_{v}(X_{3+i})|<c$, and this eigenspace can never again be extremal. Indeed, for $\lambda_{v}(X_{3+i})=e^{\pm c}$, we require $\lambda_{v}(X_{2+i})=\lambda_{v}(X_{1+i})=e^{\mp c}$, thus any extremal eigenspace for $T^{2}(X_{1},X_{0})$ will decay upon further iterations, and non-extremal eigenspaces cannot become extremal. In particular, $V(T^{k}(X_{1},X_{0}))$ must decrease as $k\to\infty$, contradicting that $E_{c}$ is invariant with fixed value function $V(E)=c$ for all $E\in E_{c}$.
In the noisy setting, Lemma 20 verifies that the recurrence remains compact for $\alpha$ sufficiently small. Hence, we can apply [Mischaikow et al., 1995, Theorem 1.8] to verify that the perturbed recurrence converges to the same limit sets as the autonomous recurrence, i.e. $I$ is the unique limit point.
@@
As in the previous proof, if we initialize with a pair in $E_{c}$, then for any $i\geq 2$, and all extremal eigenpairs $(\lambda_{\ell},v_{\ell})$, it must be that for both j=i1,i2j=i-1,i-2, the matrices $X_{j}$ have matching eigenpairs $(1/\lambda_{\ell},Jv_{\ell})$. Repeating this argument for $X_{i+1}$, for any extremal eigenpairs $(\nu_{\ell},u_{\ell})$ it follows that $X_{i},X_{i-1}$ have matching eigenpairs $(1/\nu_{\ell},Ju_{\ell})$. Combining this information, for each extremal eigenpair $(\nu_{\ell},u_{\ell})$ of $X_{i+1}$, $(1/\nu_{\ell},Ju_{\ell})$ must be an eigenpair of both $X_{i}$ and $X_{i-1}$. As $(\lambda_{\ell},v_{\ell})=(1/\nu_{\ell},Ju_{\ell})$ is an extremal eiegenpair of $X_{i}$, it must also be that $(\nu_{\ell},u_{\ell})$ is an eigenpair of $X_{i-1}$. Hence the matrix $X_{i-1}$ has both $(1/\nu_{\ell},Ju_{\ell})$ and $(\nu_{\ell},u_{\ell})$ as extremal eigenpairs, and for them to pass forward in the recurrence they must be shared by $X_{i-1}.$ Thus we see that for all $m\geq i$, $(1/\nu_{\ell},Ju_{\ell})$ and $(\nu_{\ell},u_{\ell})$ are fixed eigenvectors of $X_{m}$.
Thus we have proven that any limiting set in $E_{c}$ has fixed, cyclical extremal eigenpairs. Suppose that $X_{i}$ has $\kappa$ such limiting maximal eigenvalues. Let $\operatorname{Trunc}_{\kappa}(X)$ be the map that identifies $X$ with the $(D-\kappa)\times(D-\kappa)$ matrix formed by removing the principal $\kappa$ components. Thus, $\operatorname{Trunc}_{\kappa}T(X_{i+1},X_{i})=T(\operatorname{Trunc}_{\kappa}(X_{i},X_{i-1}))+o(1)$, and the sequence $\operatorname{Trunk}_{\kappa}(X_{i},X_{i-1})$ is asymptotically autonomous and monotone in the value function $V$. Hence [Mischaikow et al., 1995, Theorem 1.8] yields that the truncated sequence has asymptotically cyclical extremal eigenvectors.
Inductively applying this argument verifies that $X_{i}\to L$ for a fixed matrix $L$ in the noiseless case. By definition
$L=\lim_{i\to\infty}X_{i+1}=\lim_{i\to\infty}J(\eta X_{i+1}+(1-\eta)X_{i})^{-1}J=JL^{-1}J,$
hence $L\in\{X:(LJ)^{2}=I\}.$ In the noisy setting, Lemma 20 verifies that the recurrence remains compact for $\alpha$ sufficiently small. Hence, we can apply [Mischaikow et al., 1995, Theorem 1.8] to verify that the perturbed recurrence converges to the same limit sets as the autonomous recurrence.
We now study the structure of such $L$. As we demonstrated, every eigenpair $(\lambda,v)$ of $L$ must have matching eigenpair $(1/\lambda,Jv)$. Let $v_{1},\dots,v_{\kappa},Jv_{1},\dots,Jv_{\kappa},u_{1},\dots,u_{D-2\kappa}$, where $Ju_{i}=\pm u_{i}$ are invariant under $J$. For each eigenvector $u_{i}$, $\lambda=1$ is forced, thus distinct eigenvalues can only come from the 2-cycles. Displaying $J$ in this basis, we have
@@
Lemma 24 (Second Order Convergence).
Let $r\in(0,1)$, $\|R_{i}\|\leq C\alpha r^{i}$, $X_{0},X_{1}$ commute with $J$. Then, for any $\varepsilon>0$, there exists $\delta>0$ such that $X_{i}\to L_{\alpha}$, $\|L_{\alpha}-I\|<\varepsilon$.
@@
That $X_{i}$ has a limit $L_{\alpha}$ for $\alpha$ small follows immediately from Lemma 23. Thus it suffices to verify that $\lim_{i\to\infty}X_{i}=I$ for the unperturbed sequence. Indeed, if for $i$ large enough, $\|X_{i}-I\|<\varepsilon,\|X_{i-1}-I\|<\varepsilon$, because $X_{i}$ is polynomial in $X_{0},X_{1},R_{j}$, $j=1,\dots,i-1$, it follows that it is smooth for $\alpha$ small, thus incurring additional $O(\alpha)$ error. In combination with Lemma 20, this verifies the claim.
@@ B.3.6 Full Rank Hessian
Our main proof strategy is to reduce the Local EGOP iterations to the form of Proposition 2, so that we can guarantee stable first and second order convergence rates by Lemmas 19 and 24. A straightforward change of variables yields the following. Let $H$ have diagonalization $H=U\Lambda U^{T}$.
@@
$X_{i}:=W^{T}\Sigma_{i}W,\qquad\tilde{g}:=JW^{-1}g,$
@@
Then the recurrence for $X_{i}$ depends only on $J$:
$X_{i+1}=t_{i+1}J\Big($ $\tilde{g}\tilde{g}^{T}+\beta X_{i}+(1-\beta)X_{i-1}$
$+\beta[\tilde{g}\tilde{T}(X_{i})^{T}+\tilde{T}(X_{i})\tilde{g}^{T}+\tilde{R}(X_{i})]$
$+(1-\beta)[\tilde{g}\tilde{T}(X_{i-1})^{T}+\tilde{T}(X_{i-1})\tilde{g}^{T}+\tilde{R}(X_{i-1})]\Big)^{-1}J$
@@
$\alpha_{i}:=u^{T}\Sigma_{i}u,\quad\alpha_{i}^{\prime}:=u^{T}H\Sigma_{i}Hu,\quad\eta^{2}:=\|g\|^{2},\quad b_{i}:=\bar{Q}\Sigma_{i}u,\quad b_{i}^{\prime}:=\bar{Q}H\Sigma_{i}Hu,$
$C_{i}:=\bar{Q}\Sigma_{i}\bar{Q},\quad C_{i}^{\prime}:=\bar{Q}H\Sigma_{i}H\bar{Q},\quad R_{b,i}:=\eta\bar{Q}T(\Sigma_{i})+\bar{Q}R(\Sigma_{i})u,$
$\bar{R}_{\alpha,i}:=\beta[2\eta u^{T}T(\Sigma_{i})+u^{T}R(\Sigma_{i})u]+(1-\beta)[2\eta u^{T}T(\Sigma_{i-1})+u^{T}R(\Sigma_{i-1})u],$
$\bar{b}_{i}=\beta[b_{i}^{\prime}+R_{b,i}]+(1-\beta)[b_{i-1}^{\prime}+R_{b,i-1}],\qquad\bar{R}_{C,i}:=\beta\bar{Q}R(\Sigma_{i})\bar{Q}+(1-\beta)\bar{Q}R(\Sigma_{i-1})\bar{Q},$
$\bar{S}_{i}:=\beta H_{\bar{Q}}C_{i}H_{\bar{Q}}+(1-\beta)H_{\bar{Q}}C_{i-1}H_{\bar{Q}}+\bar{R}_{C,i}-\frac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}\bar{b}_{i}^{T}.$
@@
$\begin{bmatrix}\alpha_{i+1}&b_{i+1}^{T}\\ b_{i+1}&C_{i+1}\end{bmatrix}=t_{i+1}$
$\begin{bmatrix}\displaystyle\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}+\left(\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}\right)^{2}\bar{b}_{i}^{T}\bar{S}_{i}^{-1}\bar{b}_{i}&\displaystyle-\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}\bar{b}_{i}^{T}\bar{S}_{i}^{-1}\\ \displaystyle-\bar{S}_{i}^{-1}\bar{b}_{i}\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}&\displaystyle\Big(\beta C_{i}^{\prime}+(1-\beta)C_{i-1}^{\prime}+\bar{R}_{C,i}-\frac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}\bar{b}_{i}^{T}\Big)^{-1}\!\end{bmatrix}$.
@@
This is the Schur complement with pivot on the top-left scalar applied to $M_{i}:=[\beta\mathcal{L}(\mu_{i})+(1-\beta)\mathcal{L}(\mu_{i-1})]^{-1}$. ∎
@@
$\alpha_{i}=\Theta(t_{i}),\quad b_{i}=O(\sqrt{t_{i}}),\quad C_{i}=\Theta(\sqrt{t_{i}}),\quad\bar{R}_{C,i}=O(t_{i}),\quad\bar{R}_{b,i}=O(\sqrt{t_{i}}),\quad\bar{R}_{\alpha,i}=O(\sqrt{t_{i}}).$
We argue by strong induction, with the base case being covered in Corollary 3. We argue one block at a time. First note that $\bar{b}_{i}^{T}\bar{S}_{i}^{-1}\bar{b}_{i}=O(\sqrt{t_{i}})$, hence
$\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}+\left(\frac{1}{\eta^{2}+\alpha_{i}^{\prime}+\bar{R}_{\alpha,i}}\right)^{2}\bar{b}_{i}^{T}\bar{S}_{i}^{-1}\bar{b}_{i}=\frac{1}{\eta^{2}}+O(\sqrt{t_{i}})\Longrightarrow\alpha_{i+1}=\frac{t_{i+1}}{\eta^{2}}+O\left(t_{i+1}^{3/2}\right).$
Thus, taking $\alpha$ sufficiently small, we get a uniform $O(t_{i})$ bound by Lemma 19 and Proposition 3.
@@
$C_{i+1}=t_{i+1}\Big(\beta H_{\bar{Q}}C_{i}H_{\bar{Q}}+(1-\beta)H_{\bar{Q}}C_{i-1}H_{\bar{Q}}+\Big[\bar{Q}H\bar{\pi}(\beta\Sigma_{i}+(1-\beta)\Sigma_{i-1})\bar{\pi}H\bar{Q}+\bar{R}_{C,i}-\tfrac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}\bar{b}_{i}^{T}\Big]\Big)^{-1}.$
@@
$\Big[\bar{Q}H\bar{\pi}(\beta\Sigma_{i}+(1-\beta)\Sigma_{i-1})\bar{\pi}H\bar{Q}+\bar{R}_{C,i}-\tfrac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}\bar{b}_{i}^{T}\Big]/\sqrt{t_{i}}\leq C\sqrt{t_{i}}=C\alpha r^{i}$
is uniformly summable in $\alpha$, we can apply Propositions 2 and 3, and Lemma 23 to verify $C_{i+1}=\Theta(\sqrt{t_{i}})$.
Moving on to the off-diagonal, $\frac{1}{\eta^{2}+\alpha_{i}+\bar{R}_{\alpha,i}}\bar{b}_{i}^{T}\bar{S}_{i}^{-1}=O(1)$, thus upon multiplying by $t_{i+1}$, this is $O(t_{i})$, whereas the subsequent remainder terms $R_{b,i+1}=\Theta(\|\Sigma_{i+1}\|)=O(\sqrt{t_{i+1}})$. Thus, the contribution is negligible at the following iteration, and the uniform rate follows.
We can immediately plug these rates into our risk bounds to yield $\sqrt{\det\Sigma_{i}}=\Theta(t_{i}^{(D+1)/4})$, $W(\mu_{t})=O(t_{i})$, thus yielding the desired rate. ∎

### v3

@@ B.3.5 Matrix Recurrence
For this to occur, it must be that both $X_{2},X_{1}$ have eigenvalue $e^{-c}$ with a common eigenvector $v$. As the sequence is rational in $X_{2},X_{1}$, $v$ will be an eigenvector for all subsequent matrices $X_{t}$. Notice $\log\lambda_{v}(X_{4})=-\log(\eta e^{c}+(1-\eta)e^{-c})\in(-c,c)$, hence $\sup_{t>0}|\log\lambda_{v}(X_{3+t})|<c$, and this eigenspace can never again be extremal. Indeed, for $\lambda_{v}(X_{3+t})=e^{\pm c}$, we require $\lambda_{v}(X_{2+t})=\lambda_{v}(X_{1+t})=e^{\mp c}$, thus any extremal eigenspace for $T^{2}(X_{1},X_{0})$ will decay upon further iterations, and non-extremal eigenspaces cannot become extremal. In particular, $V(T^{k}(X_{1},X_{0}))$ must decrease as $k\to\infty$, contradicting that $E_{c}$ is invariant with fixed value function $V(E)=c$ for all $E\in E_{c}$.
In the noisy setting, Lemma 20 verifies that the recurrence remains compact for $\alpha$ sufficiently small. Hence, we can apply Mischaikow et al. (1995, Theorem 1.8) to verify that the perturbed recurrence converges to the same limit sets as the autonomous recurrence, i.e. II is the unique limit point. ∎
@@
As in the previous proof, if we initialize with a pair in $E_{c}$, then for any $t\geq 2$, and all extremal eigenpairs $(\lambda_{\ell},v_{\ell})$, it must be that for both j=t1,t2j=t-1,t-2, the matrices $X_{j}$ have matching eigenpairs $(1/\lambda_{\ell},Jv_{\ell})$. Repeating this argument for $X_{t+1}$, for any extremal eigenpairs $(\nu_{\ell},u_{\ell})$ it follows that $X_{t},X_{t-1}$ have matching eigenpairs $(1/\nu_{\ell},Ju_{\ell})$. Combining this information, for each extremal eigenpair $(\nu_{\ell},u_{\ell})$ of $X_{t+1}$, $(1/\nu_{\ell},Ju_{\ell})$ must be an eigenpair of both $X_{t}$ and $X_{t-1}$. As $(\lambda_{\ell},v_{\ell})=(1/\nu_{\ell},Ju_{\ell})$ is an extremal eiegenpair of $X_{t}$, it must also be that $(\nu_{\ell},u_{\ell})$ is an eigenpair of $X_{t-1}$. Hence the matrix $X_{t-1}$ has both $(1/\nu_{\ell},Ju_{\ell})$ and $(\nu_{\ell},u_{\ell})$ as extremal eigenpairs, and for them to pass forward in the recurrence they must be shared by $X_{t-1}.$ Thus we see that for all $m\geq t$, $(1/\nu_{\ell},Ju_{\ell})$ and $(\nu_{\ell},u_{\ell})$ are fixed eigenvectors of $X_{m}$.
Thus we have proven that any limiting set in $E_{c}$ has fixed, cyclical extremal eigenpairs. Suppose that $X_{t}$ has $\kappa$ such limiting maximal eigenvalues. Let $\operatorname{Trunc}_{\kappa}(X)$ be the map that identifies $X$ with the $(D-\kappa)\times(D-\kappa)$ matrix formed by removing the principal $\kappa$ components. Thus, $\operatorname{Trunc}_{\kappa}T(X_{t+1},X_{t})=T(\operatorname{Trunc}_{\kappa}(X_{t},X_{t-1}))+o(1)$, and the sequence $\operatorname{Trunk}_{\kappa}(X_{t},X_{t-1})$ is asymptotically autonomous and monotone in the value function $V$. Hence Mischaikow et al. (1995, Theorem 1.8) yields that the truncated sequence has asymptotically cyclical extremal eigenvectors.
Inductively applying this argument verifies that $X_{t}\to L$ for a fixed matrix $L$ in the noiseless case. By definition
$L=\lim_{t\to\infty}X_{t+1}=\lim_{t\to\infty}J(\eta X_{t+1}+(1-\eta)X_{t})^{-1}J=JL^{-1}J,$
hence $L\in\{X:(LJ)^{2}=I\}.$ In the perturbed setting, Lemma 20 verifies that the recurrence remains compact for $\alpha$ sufficiently small. Hence, we can once again apply Mischaikow et al. (1995, Theorem 1.8) to verify that the recurrence converges to the same limit sets as the autonomous recurrence.
We now study the structure of such $L$. As we demonstrated, every eigenpair $(\lambda,v)$ of $L$ must have matching eigenpair $(1/\lambda,Jv)$. Let $v_{1},\dots,v_{\kappa},Jv_{1},\dots,Jv_{\kappa},u_{1},\dots,u_{D-2\kappa}$, where $Ju_{i}=\pm u_{i}$ are invariant under $J$. Displaying $J$ in this basis, we have
@@
Lemma 24 (Commuting Seeds).
Let $r\in(0,1)$, $\|R_{t}\|\leq C\alpha r^{t}$, $X_{0},X_{1}$ commute with $J$. Then, for any $\varepsilon>0$, there exists $\delta>0$ such that $X_{t}\to L_{\alpha}$, $\|L_{\alpha}-I\|<\varepsilon$.
@@
That $X_{t}$ has a limit $L_{\alpha}$ for $\alpha$ small follows immediately from Lemma 23. Thus it suffices to verify that $\lim_{t\to\infty}X_{i}=I$ for the unperturbed sequence. Indeed, if for $t$ large enough, $\|X_{t}-I\|<\varepsilon,\|X_{t-1}-I\|<\varepsilon$, because $X_{t}$ is rational in $X_{0},X_{1},R_{j}$, $j=1,\dots,i-1$, it follows that it is smooth for $\alpha$ small, thus incurring additional $O(\alpha)$ error. In combination with Lemma 20, this verifies the claim.
@@ B.3.6 Full Rank Hessian
Our main proof strategy is to reduce the Local EGOP iterations to the form of Proposition 2, so that we can guarantee stable first and second order convergence rates by Lemmas 19 and 23. Let $H$ have diagonalization $H=U\Lambda U^{T}$. A straightforward change of variables yields the following.
@@
$X_{t}:=W^{T}\Sigma_{t}W,\qquad\tilde{g}:=JW^{-1}g,$
@@
Then the recurrence for $X_{t}$ depends only on $J$:
$X_{t+1}=t_{t+1}J\Big($ $\tilde{g}\tilde{g}^{T}+\beta X_{t}+(1-\beta)X_{t-1}$
$+\beta[\tilde{g}\tilde{T}(X_{t})^{T}+\tilde{T}(X_{t})\tilde{g}^{T}+\tilde{R}(X_{t})]$
$+(1-\beta)[\tilde{g}\tilde{T}(X_{t-1})^{T}+\tilde{T}(X_{t-1})\tilde{g}^{T}+\tilde{R}(X_{t-1})]\Big)^{-1}J$
@@
$\alpha_{t}:=u^{T}\Sigma_{t}u,\quad\alpha_{t}^{\prime}:=u^{T}H\Sigma_{t}Hu,\quad\eta^{2}:=\|g\|^{2},\quad b_{t}:=\bar{Q}\Sigma_{t}u,\quad b_{t}^{\prime}:=\bar{Q}H\Sigma_{t}Hu,$
$C_{t}:=\bar{Q}\Sigma_{t}\bar{Q},\quad C_{t}^{\prime}:=\bar{Q}H\Sigma_{t}H\bar{Q},\quad R_{b,t}:=\eta\bar{Q}T(\Sigma_{t})+\bar{Q}R(\Sigma_{t})u,$
$\bar{R}_{\alpha,t}:=\beta[2\eta u^{T}T(\Sigma_{t})+u^{T}R(\Sigma_{t})u]+(1-\beta)[2\eta u^{T}T(\Sigma_{t-1})+u^{T}R(\Sigma_{t-1})u],$
$\bar{b}_{t}=\beta[b_{t}^{\prime}+R_{b,t}]+(1-\beta)[b_{t-1}^{\prime}+R_{b,t-1}],\qquad\bar{R}_{C,t}:=\beta\bar{Q}R(\Sigma_{t})\bar{Q}+(1-\beta)\bar{Q}R(\Sigma_{t-1})\bar{Q},$
$\bar{S}_{t}:=\beta H_{\bar{Q}}C_{t}H_{\bar{Q}}+(1-\beta)H_{\bar{Q}}C_{t-1}H_{\bar{Q}}+\bar{R}_{C,t}-\frac{1}{\eta^{2}+\alpha_{t}+\bar{R}_{\alpha,t}}\bar{b}_{t}\bar{b}_{t}^{T}.$
@@
$\begin{bmatrix}\alpha_{t+1}&b_{t+1}^{T}\\ b_{t+1}&C_{t+1}\end{bmatrix}=t_{t+1}$
$\begin{bmatrix}\displaystyle\frac{1}{\eta^{2}+\alpha_{t}^{\prime}+\bar{R}_{\alpha,t}}+\left(\frac{1}{\eta^{2}+\alpha_{t}^{\prime}+\bar{R}_{\alpha,t}}\right)^{2}\bar{b}_{t}^{T}\bar{S}_{t}^{-1}\bar{b}_{t}&\displaystyle-\frac{1}{\eta^{2}+\alpha_{t}^{\prime}+\bar{R}_{\alpha,t}}\bar{b}_{t}^{T}\bar{S}_{t}^{-1}\\ \displaystyle-\bar{S}_{t}^{-1}\bar{b}_{t}\frac{1}{\eta^{2}+\alpha_{t}^{\prime}+\bar{R}_{\alpha,t}}&\displaystyle\Big(\beta C_{t}^{\prime}+(1-\beta)C_{t-1}^{\prime}+\bar{R}_{C,t}-\frac{1}{\eta^{2}+\alpha_{t}+\bar{R}_{\alpha,t}}\bar{b}_{t}\bar{b}_{t}^{T}\Big)^{-1}\!\end{bmatrix}$.
@@
This is the Schur complement with pivot on the top-left scalar applied to $M_{t}:=[\beta\mathcal{L}(\mu_{t})+(1-\beta)\mathcal{L}(\mu_{t-1})]^{-1}$. ∎
@@
$\alpha_{t}=\Theta(h_{t}),\quad b_{t}=O(\sqrt{h_{t}}),\quad C_{t}=\Theta(\sqrt{h_{t}}),\quad\bar{R}_{C,t}=O(h_{t}),\quad\bar{R}_{b,t}=O(\sqrt{h_{t}}),\quad\bar{R}_{\alpha,t}=O(\sqrt{h_{t}}).$
We argue by strong induction, with the base case being covered in Corollary 3. We argue one block at a time. First note that $\bar{b}_{t}^{T}\bar{S}_{t}^{-1}\bar{b}_{t}=O(\sqrt{h_{t}})$, hence
$\frac{1}{\eta^{2}+\alpha_{t}^{\prime}+\bar{R}_{\alpha,t}}+\left(\frac{1}{\eta^{2}+\alpha_{t}^{\prime}+\bar{R}_{\alpha,t}}\right)^{2}\bar{b}_{t}^{T}\bar{S}_{t}^{-1}\bar{b}_{t}=\frac{1}{\eta^{2}}+O(\sqrt{h_{t}})\Longrightarrow\alpha_{t+1}=\frac{h_{t+1}}{\eta^{2}}+O\left(h_{t+1}^{3/2}\right).$
Thus, taking $\alpha$ sufficiently small, we get a uniform $O(h_{t})$ bound by Lemma 19 and Proposition 3.
@@
$C_{t+1}=h_{t+1}\Big(\beta H_{\bar{Q}}C_{t}H_{\bar{Q}}+(1-\beta)H_{\bar{Q}}C_{t-1}H_{\bar{Q}}+\Big[\bar{Q}H\bar{\pi}(\beta\Sigma_{t}+(1-\beta)\Sigma_{t-1})\bar{\pi}H\bar{Q}+\bar{R}_{C,t}-\tfrac{1}{\eta^{2}+\alpha_{t}+\bar{R}_{\alpha,t}}\bar{b}_{t}\bar{b}_{t}^{T}\Big]\Big)^{-1}.$
@@
$\Big[\bar{Q}H\bar{\pi}(\beta\Sigma_{t}+(1-\beta)\Sigma_{t-1})\bar{\pi}H\bar{Q}+\bar{R}_{C,t}-\tfrac{1}{\eta^{2}+\alpha_{t}+\bar{R}_{\alpha,t}}\bar{b}_{t}\bar{b}_{t}^{T}\Big]/\sqrt{h_{t}}\leq C\sqrt{h_{t}}=C\alpha r^{t}$
is uniformly summable in $\alpha$, we can apply Propositions 2 and 3, and Lemma 23 to verify $C_{t+1}=\Theta(\sqrt{h_{t}})$.
Moving on to the off-diagonal, $\frac{1}{\eta^{2}+\alpha_{t}+\bar{R}_{\alpha,t}}\bar{b}_{t}^{T}\bar{S}_{t}^{-1}=O(1)$, thus upon multiplying by $h_{t+1}$, this is $O(h_{t})$, whereas the subsequent remainder terms $R_{b,t+1}=\Theta(\|\Sigma_{t+1}\|)=O(\sqrt{h_{t+1}})$. Thus, the contribution is negligible at the following iteration, and the uniform rate follows.
We can immediately plug these rates into our risk bounds to yield $\sqrt{\det\Sigma_{t}}=\Theta(h_{t}^{(D+1)/4})$, $W(\mu_{t})=O(h_{t})$, thus yielding the desired rate. ∎